OpenStudy (svnnynight9):

PLEASE HELP!!! Given the piecewise function shown below, select all of the statements that are true.(i will add the equation soon) A) f(-1)=2 B)f(-2)=0 C)f(1)=0 D)f(4)=7

OpenStudy (svnnynight9):

\[f(x)=\left\{ -x+1,x <0 \right\}\left\{ -2,x=0 \right\}\left\{ x^{2} -1,x >0\right\}\]

satellite73 (satellite73):

\(-1<0\) right?

OpenStudy (svnnynight9):

yea

satellite73 (satellite73):

meaning to find \(f(-1)\) you use the first equation

OpenStudy (svnnynight9):

so i just plug in -1 for the x?

satellite73 (satellite73):

yes

OpenStudy (svnnynight9):

so A is one of the correct statements?

satellite73 (satellite73):

yes

OpenStudy (svnnynight9):

then for f(-2)=0 i plug it in the same equation?

satellite73 (satellite73):

for \(f(-2)\) you plug \(-2\) in to the same equation, since \(-2<0\)

OpenStudy (svnnynight9):

okay and i get 3 when i plug it in

satellite73 (satellite73):

yeah so second one is not true

OpenStudy (svnnynight9):

okay so then i plug in the f(1)=0 in the first equation and get 0 so does that mean its correct??

satellite73 (satellite73):

\(1>0\) so you use the last equation

satellite73 (satellite73):

so yes, it is correct

OpenStudy (svnnynight9):

how do you know which equation to use?? @Satellite73

satellite73 (satellite73):

look to see which is true about your input

satellite73 (satellite73):

if it is less then zero use \(-x+1\) if it is 0, use \(-2\) if it is greater than zero use \(x^2-1\)

OpenStudy (svnnynight9):

so for f(4)=7 i would plug it in to the last equation

satellite73 (satellite73):

\[f(x) = \left\{\begin{array}{rcc} -x+1 & \text{if} & x <0 \\ -2 & \text{if}&x=0\\ x^2-1& \text{if} & x>0 \end{array} \right. \] is how it should read

satellite73 (satellite73):

ok i see why you would be confused the question is goofy they should just ask "what is \(f(7)\) ?"

satellite73 (satellite73):

i mean "what is \(f(4)\)?"

satellite73 (satellite73):

even i got confused since \(4>0\) use the last equation

OpenStudy (svnnynight9):

okay and i get 15

OpenStudy (svnnynight9):

so would it be false? but i still don't really get how its true or false

OpenStudy (svnnynight9):

so only A and C are correct????

satellite73 (satellite73):

A is

satellite73 (satellite73):

and C

OpenStudy (svnnynight9):

thank you so much